I keep having issues with these type of problems. Please help!
x+5y= 9
3x- y=-5
Thank you!!

Answer by RAY100(1637)

1) x+5y = 9,,,,,or x=9-5y
2) 3x -y =-5
.
using substitution,,,subst (1) into (2)
.
3(9-5y) -y = -5
27 -15y -y = -5
-16y = -32
y= 2
.
subst in (1),,,x= 9 -5(2) = -1
.
check,,,(1) (-1) +5(2) = 9,,,,ok
(2) 3(-1) -(2) = -5,,,,,ok
,
,
another technique is to multiply equations until adding or substracting them eliminates one variable
.
Multiply (2) by 5,,,(2) 15x-5y = -25
(1) x+5y=9
.
adding (1) + (2),,,,16x = -16,,,,x=-1
.
subst in (1) (-1) +5y = 9,,,5y = 10,,,y=2
.
the other technique is to graph both and note intersection.
.
Have a good day